A force F (vector)=(6i-2j)N acts on a particle undergoes a displacement delta r (Vector)=(3i+J)m. Find (a) the work done by the force on the particle and (b) the angle between F vector and r vector.
1
Expert's answer
2020-10-12T07:44:50-0400

a) The work is given by the following dot product:


"A = \mathbf{F}\cdot \mathbf{r}"

where "\mathbf{F} = 6i - 2j" and "\mathbf{r} = 3i + j". Thus, obtain:


"A = \mathbf{F}\cdot \mathbf{r} = (6i - 2j)\cdot (3i + j) = (6\cdot 3) + (-2)\cdot 1 = 16J"

b) The angle between two vectors is:


"\cos(\hat{\mathbf{F}\mathbf{r}}) = \dfrac{\mathbf{F}\cdot \mathbf{r}}{F\cdot r}"

where "F = \sqrt{\mathbf{F}\cdot \mathbf{F}} = \sqrt{6^2 + 2^2} = 2\sqrt{10}" is the lenght of vector "\mathbf{F}" and "r = \sqrt{\mathbf{r}\cdot \mathbf{r}} = \sqrt{3^2 + 1^2} = \sqrt{10}" is the length of vector "\mathbf{r}". Thus, obtain:


"\cos(\hat{\mathbf{F}\mathbf{r}}) = \dfrac{\mathbf{F}\cdot \mathbf{r}}{F\cdot r} = 
 \dfrac{16}{2\sqrt{10}\cdot\sqrt{10}} = 0.8"

The angle is:


"\alpha = \arccos(0.8) \approx 36.9\degree"

Answer. a) "16J", b) "36.9\degree".